OpenStudy (anonymous):

Need help solving for b. T=s(ab-bc)+r^2

OpenStudy (anonymous):

t-r^2=sab-sbc then factor out b...........t-r^2=b(sa-sc) then divide...... (t-r^2)/(sa-sc)=b

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

Yes, thank you so much!

OpenStudy (anonymous):

Ur welcome